Luzadis Company makes furniture using the latest automated technology. The company uses a job-order costing system and applies manufacturing overhead cost to products on the basis of machine-hours. The predetermined overhead rate was based on a cost formula that estimates $680,000 of total manufacturing overhead for an estimated activity level of 85,000 machine-hours During the year, a large quantity of furniture on the market resulted in cutting back production and a buildup of furniture in the companys warehouse. The companys cost records revealed the following actual cost and operating data for the year. Machine-hours Manufacturing overhead cost Inventories at year-end: 76,000 Raw materials Work in process (includes overhead applied of $36,480) Finished goods (includes overhead applied of $91,200) $ 637,000 $ 20,000 $115,800 $ 289,500 $1,524,700 Cost of goods sold (includes overhead applied of $480,320) Required: 1. Compute the underapplied or overapplied overhead 2. Assume that the company closes any underapplied or overapplied overhead to Cost of Goods Sold. Prepare the appropriate journal entry 3. Assume that the company allocates any underapplied or overapplied overhead proportionally to Work in Process, Finished Goods, and Cost of Goods Sold. Prepare the appropriate journal entry 4. How much higher or lower will net operating income be if the underapplied or overapplied overhead is allocated to Work in Process, Finished Goods, and Cost of Goods Sold rather than being closed to Cost of Goods Sold? Complete this question by entering your answers in the tabs below. Required Required Required Required 1 2 3 4 Compute the underapplied or overapplied overhead < Required 1 Required 2Required Required Required Required 1 2 3 4 Assume that the company allocates any underapplied or overapplied overhead proportionally to Work in Process, Finished Goods, and Cost of Goods Sold. Prepare the appropriate journal entry. (If no entry is required for a transaction/event, select No journal entry required in the first account field.) Show less View transaction list Journal entry worksheet Record the allocation of the underapplied/overapplied overhead to various accounts. Note: Enter debits before credits. Event General Journal Debit Credit Record entry View general journal Clear entryRequired Required Required Required Required 2 overhead is allocated to Work in Process, Finished Goods, and Cost of Goods Sold rather 2 3 4 gher or lower will net operating income be if the underapplied or overapplied than being closed to Cost of Goods Sold? Show less Net operating income will be overhead is allocated rather than closed entirely to cost of goods sold if the K Required 3 Required 4
Answer #1
Estimated manufacturing overhead 680000
Divide by Estimated machine hours 85000
Overhead rate 8
X Actual machine hours 76000
Overhead applied 608000
1
Actual manufacturing overhead 637000
Less: Overhead applied 608000
Underapplied overhead 29000
2
Debit Credit
Cost of goods sold 29000
          Manufacturing Overhead 29000
3
Work in Process 1740 =29000*36480/608000
Finished goods 4350 =29000*91200/608000
Cost of goods sold 22910 =29000*480320/608000
          Manufacturing Overhead 29000
4
Difference in net operating income 6090 =29000-22910
Net operating income will be $6090 higher, if the underapplied overhead is allocated rather than closed
entirely to cost of goods sold